# Sparse Narrow-Band Topology Optimization for Large-Scale Thermal-Fluid Applications

### v1

@@ Abstract
We propose a fluid-based topology-optimization methodology for convective heat-transfer problems that can manage an extensive number of design variables, enabling the fine geometric features required for the next generation of heat-exchangers design. Building on the classical Borrvall–Petersson formulation for the Stokes flow, we propose a narrow-band optimization algorithm that focuses the computational effort to the fluid–solid interface where it is most needed. To address the high cost of repeated forward and adjoint analyses, we utilize a flow solver specifically optimized for high-resolution voxel grids. The solver reduces memory usage and computational time by removing solid voxels from the analyses and directly imposing the no-slip boundary conditions on the fluid–solid interface. It also employs an efficient preconditioner built upon the established Algebraic Multigrid method that ensures fast and reliable convergence for intricate flow configurations. The discretization employs a staggered-grid finite difference scheme (Marker-and-Cell) for the Stokes-Brinkman model and an upwind finite difference scheme for the heat convection–diffusion equation, ensuring stability at high Péclet numbers. We demonstrate the method on multiple examples, including the optimization of a two-fluid heat exchanger at $Pe=10^{4}$ on a $370^{3}$ grid comprising $5\times 10^{7}$ design variables using only a single desktop workstation. The framework shows considerable promise for advancing large-scale thermal-fluid applications and constitutes an important step toward a full conjugate-heat-transfer design methodology for high-Reynolds-number Navier–Stokes flows.
@@ 1 Introduction
Topology optimization has discovered non-intuitive configurations that achieve high performance under complex physical constraints. Fluid-based topology optimization for thermal management applications has been applied to diverse areas ranging from microelectronics [1, 2] and aerospace [3] to energy storage systems including with phase change materials [4, 5]. Early work in structural topology optimization [6, 7, 8] laid the foundation for extending these methods to fluid mechanics, where researchers have addressed both laminar and turbulent flows governed by the Stokes and Navier–Stokes equations [9, 10, 11, 12]. The review from Alexandersen and Andreasen [13] summarizes the advances in the field and highlights continued research needs on topics such as computational efficiency, manufacturing constraints, and multi-objective formulations. However, the prohibitive computational cost of large-scale, three-dimensional problems remains a challenge.
@@
One of the earliest density-based methods for fluid problems was proposed by Borrvall and Petersson [9], who first suggested using the Brinkman penalization term to distinguish between “fluid” and “solid” regions during topology optimization with Stokes flow. The flow is penalized in the solid domain, typically by modeling it as a porous material with a very low permeability. Gersborg-Hansen et al. [14] later demonstrated how this approach can be extended to steady laminar Navier-Stokes flow at a moderate Reynolds number. Olesen et al. [15] investigated another integer-type extension of the Borvall’s approach to the Navier-Stokes flow. Alongside purely fluid problems, several works have studied topology optimization for conjugate heat transfer. For example, Dede [16] investigated multiphysics topology optimization of heat-transfer-driven fluid systems, while Alexandersen et al. [17] explored large-scale topology optimization of three-dimensional heat sinks with natural convection, handling up to $65.5\times 10^{6}$ design variables in a quarter domain and a total of $330\times 10^{6}$ degrees of freedom. Fawaz et al. [18], reviewing and categorizing over 200 publications on topology optimization of heat exchangers, identified the substantial computational cost of repeatedly solving flow equations as one of the primary barriers to its industrial adoption. They noticed that the resulting topologies often become highly complex, adding challenges to robust convergence of the flow solver and manufacturability. Another branch of research studied heat exchangers with multiple fluids. Notable examples include Høghøj et al. [19], who applied a density-based optimization to handle two fluids within a single domain, and Feppon et al. [20], who used a body-fitted meshing strategy for high-resolution fluid–solid boundaries. Kobayashi et al. [21] also explored conjugate heat transfer in two-fluid thermal systems using a single design variable field to represent the two fluids and the solid domain.
@@
Exhaustive reviews indicate that density-based methods using Brinkman penalization are the common approach for fluid-based topology optimization, accounting for more than 80% of the published research in this field [13, 22, 18]. The advantages of the Brinkman penalization approach include the simplicity of structured-grid discretizations, eliminating the need for discontinuity treatments at the interface. Furthermore, they allow for well-established approach for computing sensitivities. Nonetheless, several challenges remain. First, the Brinkman term only approximates the no-slip boundary condition at the fluid–solid interface. Consequently, the penalization factor must be large enough to ensure negligible flow velocity within the solid domain. However, it is known that the convergence of iterative solvers for the Stokes–Brinkman equations is highly sensitive to the penalization parameter [13]. Although various preconditioners and numerical methods have been proposed [23, 24, 25, 26, 27], a robust numerical solution strategy for heterogeneous permeability coefficient with large jumps has yet to be established. The numerical solution strategy for the Stokes–Brinkman equations continues to be an active area of research. Second, the density-based methods typically pose an optimization problem over the entire design domain, which means the velocity and pressure fields are also present in the solid region. In three-dimensional settings, this can quickly lead to a prohibitively large number of degrees of freedom. Even a moderate $400^{3}$ grid results in $64\times 10^{6}$ design variables, with roughly half a billion total unknowns (including three velocity components, pressure, and temperature), inevitably requiring a distributed memory computing resources. In addition, density-based optimization methods often employ parameter continuation (or homotopy) schemes [28, 29] for improved convergence, further increasing the computational cost due to an increase in the number of iterations.
To address these challenges, a number of strategies have been proposed. One well-known approach is to use body-fitted meshes with a level-set representation of the geometry (see, e.g., [20]). Unstructured meshes can reduce memory usage and computational effort when solving physics equations; however, frequent remeshing at every design modification becomes a serious bottleneck in three-dimensional applications. Another strategy to avoid Brinkman penalization is to use surface-capturing level-set methods, where the interface is explicitly defined by cutting the uniform background grid. In these methods, specialized integration techniques are typically employed for the cut elements, or the interface boundary conditions are enforced using a Lagrange multiplier technique. Examples of surface-capturing approaches include XFEM [30, 31, 32], cutFEM [33], the moments method [34], and immersed boundary techniques [35, 36].
@@
To reduce the computation cost and alleviate the memory bottleneck of dense grid methods, Kambampati et al. [37] proposed an approach using an efficient VDB data structure [38] to represent designs in topology optimization. The VDB data structure is based on sparse grids, which enables efficient allocation of memory and computational effort without a significant overhead. An illustrative example of this sparse grid concept is the propagation of level sets, where memory and computational resources are utilized only within a narrow band around the evolving interface, avoiding recalculation across the entire computational domain [39]. We adopt this narrow-band concept in our optimization algorithm, described in detail in Section 3. Specifically, during topology optimization, design variables are dynamically activated or deactivated, ensuring that only a small subset remains active at each optimization step. Consequently, the optimization problem is transformed into a sequence of Integer Linear Programs (ILPs) solved iteratively for these active subsets using the Sequential Linear Programming (SLP) approach. This methodology is similar to the Topology Optimization of Binary Structures (TOBS) algorithm [40, 41, 42]. However, to maintain the validity of linear approximations, TOBS introduces an additional constraint limiting the design variations between two subsequent iterations, while our approach explicitly confines active subsets to a narrow band around the fluid-solid interface. This narrow-band optimization methodology allows us to leverage the sparsity advantage: Since design updates are restricted to the predefined active subset around the fluid-solid interface, solid voxels that are known to remain solid throughout optimization can be excluded from the analysis, offering significant computational acceleration. In fluid topology optimization, the fluid volume fraction may be low, in which case a significant portion of computational effort is spent unnecessarily on fictitious solid regions. This sparsity concept was proposed in fluid-based level-set topology optimization by Challis and Guest [43], who advocated eliminating fictitious solid voxels and imposing no-slip boundary conditions directly on the fluid-solid interface. By discarding degrees of freedom associated with these solids, they enabled solving large-scale three-dimensional problems using a basic iterative method without any special preconditioning. In the current study, we extend this approach to conjugate heat transfer problems and further accelerate it with an efficient numerical method. A similar idea of imposing the no-slip boundary conditions on the fluid-solid interface was employed in [44] and [42].
Outside the context of topology optimization, the sparsity concept of excluding fictitious solids is extensively applied in the field of Digital Rock Physics (DRP) [45, 46, 47]. In DRP, the primary objective is to simulate micro-scale fluid flow through porous media, deriving effective macroscopic properties from high-resolution voxel grids that correspond directly to computed-tomography images with up to $8\times 10^{9}$ voxels [48]. Like fluid-based topology optimization, DRP extensively utilizes structured voxel grids and Brinkman penalization to handle unresolved or grayscale voxels, posing a shared computational challenge in efficiently solving the Stokes– Brinkman equations. Additionally, DRP solvers are developed to perform consistently across a wide range of geometries because of the significant variability in pore structures encountered in practical applications. In topology optimization, this translates to the requirement for solvers to maintain stable and reliable performance over the entire design space. It is worth emphasizing that flow solvers developed for simple geometries often do not perform well for complicated geometries, such as those arising in porous-media applications or fluid-based topology optimization. Recognizing these parallels, in the current study we adapt an efficient flow solver originally developed for porous-media applications by Pimanov et al. [27, 49] to fluid-based topology optimization. Details on the discretization and numerical method are provided in Sections 4 and 5, respectively.
@@
Our key contributions can be summarized as follows: (i) We propose a narrow-band topology-optimization algorithm that focuses on the evolving fluid–solid interface; (ii) We integrate an efficient flow solver tailored to high-resolution voxelized grids and allowing to exclude fictitious solid voxels; and (iii) We provide a detailed implementation, including sensitivity analysis and discretization schemes for both fluid flow and heat transfer. We demonstrate the method on a suite of examples, including a pure-fluid problem, a conjugate heat-transfer case, and a two-fluid heat exchanger with a minimum-thickness constraint.
@@ 2 Governing equations
Following the classical approach of Borrvall and Petersson [9], we describe fluid flow throughout the entire domain $\Omega$ via introducing the Brinkman penalization term. Let $u(\mathbf{x})$ be the velocity field, and $p(\mathbf{x})$ the pressure field. The Stokes-Brinkman equations are given as follows:
@@ 3 Narrow-Band Topology Optimization
We solve the local subproblem (12) using the Sequential Linear Programming (SLP) approach [50]. In order to apply SLP, we first reformulate the optimization subproblem in the perturbation form. Namely, we look for a new design in the form:
@@
The proposed optimization method is related to the SLP method for level-set topology optimization proposed by Dunning and Kim [50], customized for fluid-based optimization. In comparison to the level-set optimization, our method operates directly on the density level and does not require an auxiliary evolution equation such as the Hamilton–Jacobi equation. Thus, the proposed narrow-band optimization algorithm can be seamlessly integrated into standard density-based frameworks for large-scale thermal-fluid topology optimization. The overlap and similarities between level set and density-based optimization were also pointed out by Andreasen et al. [51].
@@ 4 Sparse grid discretization
The Stokes–Brinkman problem (1) is discretized using the finite-difference method on staggered grids, also known as Marker-and-Cell scheme [52, 53, 54]. In the staggered grid arrangement, the pressure grid coincides with the background grid while the components of the velocity $U$, denoted $U^{x}$ and $U^{y}$, are defined on the cell faces. For the heat equation (3), we use the centered finite-difference scheme for discretizing the Laplacian operator $\Delta\theta$ and the upwind scheme tailored for a high–Péclet number regime [55] for discretizing the convective term $\nabla\cdot\bigl{(}{u}(\gamma)\,\theta\bigr{)}$. The degrees of freedom (DoFs) are located as follows (see Fig. 3):
@@
The discrete incompressibility condition is written at the cell centers for $P_{i,j}$ as follows:

### v2

@@ Abstract
We propose a fluid-based topology optimization methodology for convective heat-transfer problems that can manage an extensive number of design variables, enabling the fine geometric features required for the next generation of heat-exchangers design. Building on the classical Borrvall–Petersson formulation for the Stokes flow, we introduce an optimization algorithm that focuses computational effort on the fluid–solid interface, where it is most needed. To address the high cost of repeated forward and adjoint analyses and to avoid leakage through nominally solid regions, we exclude fictitious solid voxels from the analysis by imposing the no‑slip boundary conditions in the vicinity of the fluid-solid interface. In contrast to the prior approaches, the fictitious solids are also excluded from the global optimization problem via reducing it to a sequence of local narrow‑band subproblems with a variable design space. The contribution of our method is that large-scale optimization can be solved efficiently by continuous simplex method while reliably obtaining binary designs without additional filtering or projection. We demonstrate efficiency of the method on multiple examples, including the optimization of a two-fluid heat exchanger at $Pe=10^{4}$ on a $370^{3}$ grid comprising $5\times 10^{7}$ design variables using only a single desktop workstation.
@@ 1 Introduction
Topology optimization has discovered non-intuitive configurations that achieve high performance under complex physical constraints. Fluid-based topology optimization for thermal management applications has been applied to diverse areas ranging from microelectronics [zeng2019topology, zou2022topology] and aerospace [kambampati2021level] to energy storage systems, including designs with phase change materials [pizzolato2017design, pizzolato2017topology]. Early work in structural topology optimization [bendsoe1988generating, bendsoe1989optimal, sigmund2013review] laid the foundation for extending these methods to fluid mechanics, where researchers have addressed both laminar and turbulent flows governed by the Stokes and Navier–Stokes equations [borrvall2003, gersborg2005topology, dilgen2018topology, yoon2016topology]. The review by Alexandersen and Andreasen [alexandersen2020review] summarizes advances in the field and highlights continued research needs on topics such as computational efficiency, manufacturing constraints, and multi-objective formulations. However, the prohibitive computational cost of large-scale, three-dimensional problems remains a challenge.
One of the earliest density-based methods for fluid problems was proposed by Borrvall and Petersson [borrvall2003], who suggested material interpolation via a Brinkman reaction term to distinguish between “fluid” and “solid” regions during topology optimization with Stokes flow. The flow is penalized in the solid domain by modeling it as a porous material with very low permeability. Gersborg-Hansen et al. [gersborg2005] later demonstrated how this approach can be extended to steady laminar Navier–Stokes flow at moderate Reynolds number. Olesen et al. [olesen2006] implemented a high-level Navier–Stokes topology optimization framework based on the fictitious porous media approach.
Alongside purely fluid problems, several works have studied topology optimization for conjugate heat transfer. Alexandersen et al. [alexandersen2016] explored large-scale topology optimization of three-dimensional heat sinks with natural convection, handling up to $65.5\times 10^{6}$ design variables. Other representative studies on conjugate heat transfer that build on the Borrvall–Petersson framework include a convection-dominated formulation by Bruns et al. [bruns2007topology], natural convection modeled with the Boussinesq approximation by Alexandersen et al. [alexandersen2014topology], forced convection with turbulent extensions by Dilgen et al. [dilgen2018density], and applications to gas-turbine cooling configurations by Pietropaoli et al. [pietropaoli2019three].
@@
Another branch of research studies heat exchangers with multiple fluids. Notable examples include Høghøj et al. [hoghoj2020], who applied a density-based optimization to handle two fluids within a single domain, and Kobayashi et al. [kobayashi2021topology] also explored conjugate heat transfer in two-fluid thermal systems using a single design-variable field to represent the two fluids and the solid domain. Fawaz et al. [fawaz2022topology], reviewing and categorizing over two hundreds publications on topology optimization of heat exchangers, identified the substantial computational cost of repeatedly solving flow equations as one of the primary barriers to industrial adoption. They noted that the resulting topologies often become highly complex, adding challenges to robust convergence of the flow solver.
@@
Exhaustive reviews indicate that density-based methods relying on Borrvall–Petersson material interpolation are the common approach for the fluid-based topology optimization, accounting for more than 80% of the published research in this field [alexandersen2020review, dbouk2017review, fawaz2022topology]. The advantages of this approach include the simplicity of structured-grid discretizations, eliminating the need for discontinuity treatments at the interface, and the availability of well-established techniques for computing sensitivities. Nonetheless, several challenges remain. First, the Brinkman term only approximates the no-slip boundary condition at the fluid–solid interface. Consequently, for faster flows, the porous-medium surrogate may exploit “flow seepage” through nominally solid regions, creating numerically permeable obstacles that would be physically impermeable. As demonstrated and discussed extensively by Kreissl and Maute [kreissl2012levelset] in a level-set context, and observed to a lesser extent in Lin et al. [lin2015topology], such leakage paths can materially degrade the performance and may violate imposed mass-flow constraints. Consequently, the penalization factor must be large enough to ensure negligible flow velocity within the solid domain. However, it is known that the convergence of iterative solvers for the Stokes–Brinkman equations is highly sensitive to the penalization parameter [alexandersen2020review]. Although various preconditioners and numerical methods have been proposed [iliev2013numerical, vassilevski2013block, efendiev2012robust, iliev2011variational, pimanov2022workflow], a robust numerical solution strategy for heterogeneous permeability coefficients with large jumps has yet to be established; the numerical solution of Stokes–Brinkman equations continues to be an active area of research. Second, density-based methods typically pose an optimization problem over the entire design domain, which means the velocity and pressure fields are also present in the solid region. In three-dimensional settings, this can quickly lead to a prohibitively large number of degrees of freedom. Even a moderate $400^{3}$ grid results in $64\times 10^{6}$ design variables, with roughly half a billion total unknowns (including three velocity components, pressure, and temperature), inevitably requiring distributed-memory computing resources. In addition, density-based optimization methods often employ parameter continuation (or homotopy) schemes [pereira2016fluid, alexandersen2023detailed] for improved convergence, further increasing the computational cost due to the growth in the number of iterations. To overcome these difficulties, a number of strategies have been proposed. One well‑known approach is to use unstructured body‑fitted meshes with a level‑set representation of the geometry (see, e.g., [feppon2021body]). Body‑fitted discretizations naturally allow solving the governing equations only in the fluid domain; however, frequent remeshing at every design modification becomes a serious bottleneck in three‑dimensional applications. Another family of methods that only solves for the fluid regions relies on surface‑capturing discretizations. Examples of surface‑capturing discretizations applied to topology optimization include XFEM [kreissl2012levelset, jenkins2015level, hoghoj2025density, noel2023xfem], cutFEM [villanueva2017cutfem], the moments method [kambampati2023cad], and immersed‑boundary techniques [jenkins2016immersed, kubo2021level]. In topology optimization settings, these discretizations are commonly paired with a level‑set geometry description: The interface is represented implicitly and then discretized by cutting a uniform background grid. In immersed‑boundary methods, in order to impose boundary conditions directly at the interface, a localized forcing (e.g., Brinkman‑type penalization or distributed Lagrange multipliers) is introduced rather than using a heuristic material interpolation. While these methods can yield an improved accuracy near the boundaries, they typically require specialized quadrature for cut elements or ghost‑cell/extended stencils. In level‑set–based topology optimization with surface‑capturing discretizations, one typically computes an interface velocity that minimizes an objective functional while respecting the constraints and advances a Hamilton–Jacobi level‑set equation to update the geometry; the associated indicator (density) field is implied by the new interface. In our approach, we do not determine the interface position explicitly. Instead, we update the indicator field directly, restricting changes to a narrow band around the current interface. Under this restriction, we assume the updated field corresponds to a small interface displacement, i.e., the new interface lies within the band.
@@
To reduce the computational cost and alleviate the memory bottleneck of uniform‑grid methods, Kambampati et al. [kambampati2021geometry] proposed representing designs using an efficient VDB data structure [museth2013vdb]. The VDB representation is based on sparse grids, enabling memory and work to be allocated only where needed with negligible overhead. An illustrative example is level‑set propagation, where memory and computation are concentrated within a narrow band around the evolving interface, avoiding unnecessary work across the entire domain [adalsteinsson1995fast]. We adopt this narrow‑band concept in our optimization algorithm, described in detail in Section 3. Specifically, during topology optimization, design variables are dynamically activated or deactivated so that only a small subset remains active at each step. Consequently, the global problem is reduced to a sequence of subproblems solved iteratively for these active subsets using a Sequential Linear Programming (SLP) strategy. This methodology is similar to the Topology Optimization of Binary Structures (TOBS) algorithm [sivapuram2018topology, souza2021topology, picelli2022topology]. However, to maintain the validity of linear approximations, TOBS introduces an additional constraint limiting design changes between successive iterations, whereas our approach explicitly confines the active set to a narrow band around the fluid–solid interface. Moreover, in contrast to the TOBS method that relies on integer optimizers, we employ continuous‑variable simplex linear programming updates within the SLP framework enabling good scaling properties.
The narrow-band optimization methodology allows us to leverage the sparsity advantage: Since design updates are restricted to the predefined active subset around the fluid-solid interface, solid voxels that are known to remain solid throughout optimization can be excluded from the analysis, offering significant computational acceleration. In fluid topology optimization, the fluid volume fraction may be low, in which case a significant portion of computational effort is spent unnecessarily on fictitious solid regions. This sparsity concept was advocated in fluid-based level-set topology optimization by Challis and Guest [challis_guest], who discussed eliminating fictitious solid voxels and imposing no-slip boundary conditions directly on the fluid-solid interface. By discarding degrees of freedom associated with these solids, they enabled solving large-scale three-dimensional problems using a basic iterative method without any special preconditioning. In the current study, we extend this idea to conjugate heat transfer problems and further accelerate it with an efficient numerical method. Most prior topology‑optimization approaches that solve only in the fluid domain rely on explicit interface descriptions (typically level‑set methods). A drawback of such level‑set‑based approaches is that they require the derivation of shape sensitivities, which is considered more cumbersome than in the Brinkman penalization approach. Excluding fictitious voxels has been shown to be effective in structural density‑based topology optimization (see, e.g., [bruns2003element]), but transferring this idea to a Brinkman framework for fluids is not straightforward: The key challenge in using adaptive, design‑dependent boundary conditions is to ensure that introducing boundary conditions does not overly restrict the design space [behrou2019adaptive, theulings2023towards]. Behrou et al. [behrou2019adaptive] proposed a density‑based optimization method for incompressible Navier–Stokes flow that utilizes adaptive no‑slip boundary conditions. The authors exclude fictitious solid voxels using a simple threshold. However, they remove them from the finite element analysis but not from the optimization problem. As a result, they acknowledge that a discontinuity into an otherwise continuous topology optimization problem is introduced, and that the designs are likely to depend on the chosen threshold magnitude (see the discussion in Section 3.1 [behrou2019adaptive]). In our proposed method, we alleviate this difficulty by removing solids also from the optimization, via considering a sequence of optimization subproblems with a variable design space instead of a single optimization problem. In [behrou2019adaptive], the authors use a Heaviside projection to obtain binary designs. A distinctive feature of our method is that, despite using a continuous optimizer (simplex) within the standard Brinkman penalization approach, we obtain binary designs without any filtering or projection. Another density-based method that reconstructs explicit interface from densities and uses cut-cell method for computing sensitivities was proposed by Galanos et al. [galanos2024cut, galanos2025continuous]. However, the Helmholtz filter [lazarov2011filters] is utilized in their work which results in a diffuse-interface.
@@
Our key contributions can be summarized as follows: (i) We propose a narrow-band topology optimization algorithm that focuses on the evolving fluid–solid interface; the proposed algorithm allows for robust removal of fictitious solids in the classical Borrvall-Petersson formulation (ii) Despite using a continuous optimizer (simplex) within the classical Brinkman framework, we obtain binary designs without any filtering or projection. (iii) We provide a detailed implementation, including sensitivity analysis and discretization schemes for both fluid flow and heat transfer. We demonstrate the method on a suite of examples, including a pure-fluid problem, a conjugate heat-transfer case, and a two-fluid heat exchanger with a minimum-thickness constraint.
@@ 2 Governing equations
Following the classical approach of Borrvall and Petersson [borrvall2003], we describe fluid flow throughout the entire domain $\Omega$ via introducing the Brinkman penalization term. Let $u(\mathbf{x})$ be the velocity field, and $p(\mathbf{x})$ the pressure field. The Stokes-Brinkman equations are given as follows:
@@ 3 Narrow-Band Topology Optimization
We solve the local subproblem (12) using the Sequential Linear Programming (SLP) approach [dunning2015introducing]. In order to apply SLP, we first reformulate the optimization subproblem in the perturbation form. Namely, we look for a new design in the form:
@@
The proposed optimization method is related to the SLP method for level-set topology optimization proposed by Dunning and Kim [dunning2015introducing], customized for fluid-based optimization. In comparison to the level-set optimization, our method operates directly on the density level and does not require an auxiliary evolution equation such as the Hamilton–Jacobi equation. Thus, the proposed narrow-band optimization algorithm can be seamlessly integrated into standard density-based frameworks for large-scale thermal-fluid topology optimization. The overlap and similarities between level set and density-based optimization were also pointed out by Andreasen et al. [andreasen2020level].
Remark 1
The fixed‑volume constraint used in (12)–(17) (i.e., $\sum_{j\in a}(\Delta\Gamma^{0,k}_{a})_{j}=0$) is not a severe limitation of the method. Varying volume during optimization can be achieved, e.g., by an additional outer loop where the target volume is adjusted by shrinking/expanding the interface.
@@ 4 Sparse grid discretization
The Stokes–Brinkman problem (1) is discretized using the finite difference method on staggered grids, also known as Marker-and-Cell (MAC) scheme [harlow1965numerical, lebedev1964difference, griebel1998numerical]. We adapt the Stokes–Brinkman solver of Pimanov et al. [pimanov2022workflow, pimanov2024efficient] to fluid‑based topology optimization. We prefer the staggered finite difference discretization over the finite element method (FEM) because (i) equal‑order velocity–pressure FEM pairs violate the discrete inf–sup condition and thus require stabilization [brezzi2012mixed, tezduyar1991stabilized], whereas the MAC scheme satisfies a discrete inf–sup condition by construction [rui2017stability]; and (ii) for Brinkman flow with high‑contrast permeability, standard Stokes‑stable pairs (e.g., Taylor–Hood) are not robust and typically need additional stabilization or specialized preconditioners [hannukainen2011computations, efendiev2012robust, vassilevski2013block].
In the staggered grid arrangement, the pressure grid coincides with the background grid while the components of the velocity $U$, denoted $U^{x}$ and $U^{y}$, are defined on the cell faces. For the heat equation (3), we use the centered finite-difference scheme for discretizing the Laplacian operator $\Delta\theta$ and the upwind scheme tailored for a high–Péclet number regime [shyy1985study] for discretizing the convective term $\nabla\cdot\bigl({u}(\gamma)\,\theta\bigr)$. The degrees of freedom (DoFs) are located as follows (see Fig. 3):
@@
The discrete incompressibility condition is written at the cell centers, where $P_{i,j}$ is defined, as follows: